### v1

@@ Abstract
Current 3D reconstruction techniques struggle to infer unbounded scenes from a few images faithfully. Specifically, existing methods have high computational demands, require detailed pose information, and cannot reconstruct occluded regions reliably. We introduce 6Img-to-3D, an efficient, scalable transformer-based encoder-renderer method for single-shot image to 3D reconstruction. Our method outputs a 3D-consistent parameterized triplane from only six outward-facing input images for large-scale, unbounded outdoor driving scenarios. We take a step towards resolving existing shortcomings by combining contracted custom cross- and self-attention mechanisms for triplane parameterization, differentiable volume rendering, scene contraction, and image feature projection. We showcase that six surround-view vehicle images from a single timestamp without global pose information are enough to reconstruct 360 scenes during inference time, taking 395 ms. Our method allows, for example, rendering third-person images and birds-eye views. Our code is available here, and more examples can be found at our website https://6Img-to-3D.GitHub.io/.
Keywords:
Scene Representation Few-View Reconstruction 3D Reconstruction Autonomous Vehicles
Figure 1: Qualitative results of 6Img-to-3D. During inference time, 6Img-to-3D takes six surround vehicle RGB images as input and returns a parameterized triplane from which arbitrary views can be rendered.
@@ 1 Introduction
Inferring the appearance and geometry of large-scale outdoor scenes from a few camera inputs is a challenging, unsolved problem. The problem is characterized by the complexity inherent in outdoor scenes, namely a vast spatial extent, diverse object textures, and ambiguity of the 3D geometry when resolving occlusions. Robotics and autonomous driving systems both require methods that process vision-centric first-person views from a fixed set of cameras as input and derive suitable control commands from these [80]. The ability to instantaneously perform 2D-to-3D scene reconstructions would be useful within such methods and thus have a broad applicability within the robotics [2, 51, 60, 89] and autonomous driving [61, 15, 111, 8] domains. Instantaneously generating unoccluded bird’s-eye views in difficult parking or slow driving scenarios would, for example, greatly assist human drivers. For applying one-shot images to high-fidelity 3D techniques in vehicles or robots, additional factors, such as the inference speed, the quality of the inferred representation, the scalability of the approach, and the number of camera parameters required for the technique to work, are equally important. For those safety-critical domains, avoiding overly compressing the 3D scene or altering scene details by adding or omitting road obstacles is crucial. For those safety-critical domains, avoiding overly compressing the 3D scene or altering scene details by adding or omitting road obstacles is crucial.
Despite recent progress in 3D reconstruction from a few or single 2D images, many methods are either limited to single objects [27, 76, 46, 45, 96] or to generative indoor view synthesis [12]. Existing few-shot methods developed for single objects are often not easily transferable to large-scale or unbounded outdoor settings. This is because outdoor settings are more diverse due to their varying structural composition, occlusions, and vast extent. Another limitation is that many methods require a high degree of overlap between the input images, such as found in spherically arranged and inward-facing camera setups focused on a single object in the center [55, 103]. This is opposed to the outward-facing camera setup essential for autonomous driving where the camera overlap is minimal [6, 9, 74]. The camera setups are contrasted in Fig. 2. Traditional methods in autonomous driving have shown promising results with such camera setups but lack visual fidelity since they only predict semantic occupancy [28, 81, 72, 87] and commonly rely on additional sensors such as LiDAR.
@@
Given the above limitations, we present 6Img-to-3D, a novel transformer-based decoder-renderer method for single-shot image to 3D reconstruction. 6Img-to-3D is fully differentiable and end-to-end trained on a single GPU. It simultaneously takes six outward-facing images (6 Img) as inputs and generates a parameterized triplane (to 3D) from which novel scene views can be rendered. Unlike previous methods, we do not focus on single objects but instead on complex driving scenarios characterized by a high-depth complexity recorded only by sparse multi-view outward-facing RGB cameras. Furthermore, unlike other methods, we do not use a multi- or few-shot scheme where an initial reconstruction is tested and improved iteratively. Instead, we provide the 3D reconstruction in a single shot without iterative model refinement. To this end, we utilize pre-trained ResNet features, cross- and self-attention mechanisms for parameterizing the triplane, and image feature projection to condition the renderer. Our method has 61M parameters and is trained using roughly 11.4K input and 190K supervising images from 1900 scenes. We do not require depth or LiDAR information. Due to the absence of suitable datasets, including a large number of non-ego vehicle views needed during training, we rely on synthetic data from Carla [19].
@@
    Visual Fidelity and Unboundedness. We present 6Img-to-3D, an efficient, single-shot, few-image reconstruction pipeline for large-scale outdoor environments. Our method outputs a parameterized triplane of the 3D scene from which arbitrary viewpoints can be rendered. We combine triplane-based differentiable volume rendering, renderer conditioning on projected image features, self- and custom cross-attention mechanisms, and an LPIPS loss.
@@
    Scaling and Hardware Efficiency. We demonstrate the favorable scaling properties of our architecture regarding the amount of training data. Unlike other large-scale 2D-to-3D methods, 6Img-to-3D trains and runs on a single 42GB GPU, allowing easier transfer of the method to embedded systems.
@@
    Self-Supervised and Runtime. We establish that our method trained by using self-supervised learning generalizes well to novel scenes without inputting additional camera poses other than those required to define the novel views. Since our method only requires a single-forward pass during inference, obtaining a parameterized triplane is fast and only takes 395ms.
Our code is available here.
@@ 2 Related Work
3D Representations. Traditional methods for representing 3D scenes include voxels [68, 24, 91, 97, 88, 75], point clouds [20, 105, 100, 1, 106, 75], signed distance field (SDF) [84, 38] and polygon meshes [75, 32, 83, 57, 94, 48]. Recently, numerous new implicit and explicit representations have emerged for learning 3D scenes [79, 92, 23]. Implicit neural fields such as Neural Radiance Fields (NeRFs) model the surrounding appearance and geometry using a continuously queriable function. They can be updated through differentiable volumetric rendering [59, 55, 49, 77, 23] and either utilize a single [17, 55, 108] or multiple [66, 64] multilayer perceptrons (MLPs). However, these models are slow to train, data-intensive, and expensive to render since querying the implicit neural field is computationally demanding. Explicit approaches such as Plenoctrees [102], Plenoxels [101], and NSVF [44] employ trilinear interpolation within a 3D grid to represent the scene. This significantly improves rendering and optimization time but comes at the expense of limited scalability due to the curse of dimensionality when increasing the resolution or scene size. Hybrid models [73, 56] adopt an intermediary representation, storing values in a voxel-like grid. These values are then decoded using viewing direction through a shallow MLP to infer color and density. Some models further refine this process by decomposing the voxel grid into vector-matrix, using the CANDECOMP/PARAFAC (CP) decompositions [13] or into three or more orthogonal planes called triplanes or K-Planes [11, 5, 21]. This enables the computation of features for each point in space by aggregating the feature values of projected 3D points onto each plane, transitioning from an N-cube voxel grid to a 3 $\times$ N-squared plane representation.
@@
Few-Image to 3D Representation. Single-scene 3D reconstruction methods require many images to generalize towards novel views since the reconstruction problem would be underconstrained otherwise [55]. Hence, few-image to 3D representation methods incorporate additional global or local information to further regularize and constrain the problem. Global regularization methods rely on additional model priors, either via applying further model regularization [98, 58] or by utilizing pre-trained image models to provide an extra guidance signal to inform the optimization of the single-scene [18, 63, 45, 53, 69, 93, 96, 42, 62]. Local regularization methods train a cross-scene multi-perspective aggregator. Such models include PixelNeRF [103], IBRNet [85], MVSNeRF [14], VolRecon [67] and others [82, 47, 31, 40]. They retrieve image features via view projection and aggregate resulting features using MLPs or transformers to obtain novel views. We also rely on pixel feature conditioning, but unlike our method, many of the mentioned local regularization methods rely on a large overlap between input images ensured by using spherical inward-facing cameras. Another category of methods does not train the representation on a per-scene basis but instead trains a meta-network that outputs the discrete parameterized scene representation. Models such as [71, 22, 25, 3, 86] use diffusion to generate 3D representations of scenes. Many of those models use the triplane representation [71, 22, 25, 3], whereas some directly generate in the image-space [86] or can be conditioned on input images [25, 86, 3]. Other methods [11, 7] use GANs to generate 3D representations. The method that most closely resembles our work is LRM [27], which utilizes cross- and self-attention to parameterize a triplane using a few object images. LRM focuses on single objects and does not handle unbound scenes. In addition, we train on a single A40 GPU, whereas LRM is trained on 128 NVIDIA A100 GPUs and has seven times more parameters.
@@
Large-Scale 3D Scene Representations. Extending NeRFs to unbounded scenes requires spatially contracting the scene is usually done proportionally to disparity as introduced in MipNeRF-360 [4] or using an inverted sphere parametrization as in NeRF++ [109]. Architectures modeling outdoor driving scenes are BlockNeRF [78], Neural Scene Graphs [61], DisCoScene [95], NeuralField-LDM [33] and Neural ground planes [70]. Those are, however, not dealing with sparse input views. Neo360 [30] utilizes local image features to infer an image-conditional triplanar representation, whereby the model dissociates foreground from background scene parts. Unlike our method, Neo360 is trained on 8 A100 GPUs and focuses on inward-facing camera perspectives as input to the model. Due to the larger view overlap, Neo360 thus solves an easier 2D-to-3D lifting task. Large-scale scene generation methods such as Infinite Nature [43], InfiniteNature-Zero [36], PathDreamer [35], CityDreamer [90], Infinicity [39], SceneDreamer [16] and PersistentNature [10] often produce additional images in an autoregressive fashion without the ability to input camera coordinates for novel view synthesis control. In autonomous driving, 3D visual perception tasks include 3D detection and map segmentation based on multi-camera images. Established methods create a bird’s-eye-view (BEV) semantic representation of a scene using either Inverse Perspective Mapping (IPM) [65] or attention mechanisms [99, 37]. TPVFormer [29], a work that our architecture is partially based on, maps six outward-facing vehicle 2D images onto three orthogonal planes using self- and cross-deformable attention mechanisms to obtain semantic occupancy predictions. The method is, however, supervised with depth information from a LiDAR sensor and, unlike our method, only predicts semantic occupancy.
@@
Given six outward-facing images $\bm{I}_{ego}$, their camera extrinsic (relative pose) matrices $\bm{M}$ and associated camera intrinsic matrices $\bm{K}$ with appropriate dimensions, the goal is to reconstruct the surrounding 3D occupancy and appearance, that is the scene $\bm{S}$. Since the values of the scene $\bm{S}$ are not known, we use ${N}$ spherical multi-view images $\bm{I}_{sphere}$ with their associated extrinsic and extrinsic matrices during training as a supervision signal to estimate the scene $\bm{\hat{S}}$. The proposed architecture for lifting $\bm{I}_{ego}$ consists of an image-to-triplane encoder. After a forward pass through this architecture, the discrete triplane contains the approximation of the scene $\bm{\hat{S}}$. The following sections describe the different parts of the pipeline visualized in Fig. 3 and the training objectives.
@@ 3.3 Image-to-Triplane Encoder
$\textbf{CrossAttn}(\bm{T_{k}})=\text{DeformAttn}(\bm{T_{k}},\bm{p^{F}_{k}},\bm{F_{I}}),k\in\{HW,HZ,WZ\}$
@@
$\textbf{SelfAttn}(\bm{T_{k}})=\text{DeformAttn}(\bm{T_{k}},\bm{p^{T}_{k}},\mathbf{T}),k\in\{HW,HZ,WZ\}\;.$
@@ 3.5 Training Objectives
$\mathcal{L}=\mathcal{L}_{color}(\bm{C}(\bm{r}),\hat{\bm{C}}(\bm{r}))+\lambda_{t}\mathcal{L}_{TV}(\bm{T})+\lambda_{d}\mathcal{L}_{dist}(\bm{s},\bm{w})+\lambda_{l}\mathcal{L}_{\text{LPIPS}}(\bm{x},\hat{\bm{x}})\;.$ (3)
@@
$\mathcal{L}_{dist}(\bm{s},\bm{w})=\sum_{i,j}w_{i}w_{j}\bigg{|}\frac{s_{i}+s_{i+1}}{2}-\frac{s_{j}+s_{j+q}}{2}\bigg{|}+\frac{1}{3}\sum_{i}w_{i}^{2}(s_{i+1}-s_{i})\;,$ (6)
where $\bm{s}$ is a set of ray distances and $\bm{w}$ are the volume rendering weights parameterizing each ray, computed following [4].
@@ 4.1 Dataset
Since datasets with both ego and non-ego vehicle views are unavailable, we introduce a novel dataset for few-view image reconstruction in an autonomous driving setting. Concerning 3D reconstruction, other approaches also rely on synthetic data [30]. Our data is produced using the open-source CARLA Simulator [19]. Our dataset contains 2000 single-timestep complex outdoor driving scenes, each offering six outward-facing vehicle images and 100 spherical images for supervision. We also produce five multi-timestep scenes, each with 200 consecutive driving steps containing six outward-facing vehicle cameras, one bird’s-eye view, and one camera following the vehicle from behind. The multi-timestep data is only used for visualization. This results in 220K individual RGB images, each with their associated intrinsic camera matrix and pose. The dataset represents various driving scenes, vehicle types, pedestrians, and lighting conditions. To show the generalizability of our method, we use CARLA [19] Town 1, Town 2 to 7, and 10 for generating training data, resulting in 1900 training scenes, and left the 100 scenes from Town 2 for validation. We plan to make the dataset and the data generator publicly available in a separate future contribution.

### v2

@@ Abstract
Current 3D reconstruction techniques struggle to infer unbounded scenes from a few images faithfully. Most existing methods have high computational demands, require detailed pose information, and cannot reconstruct occluded regions reliably. We introduce 6Img-to-3D, a novel transformer-based encoder-renderer method for single-shot image-to-3D reconstruction. Our method outputs a 3D-consistent parameterized triplane from only six outward-facing input images for large-scale, unbounded outdoor driving scenarios. We take a step towards resolving existing shortcomings by combining contracted custom cross- and self-attention mechanisms for triplane parameterization, differentiable volume rendering, scene contraction, and image feature projection. We showcase on synthetic data that six surround-view vehicle images from a single timestamp are enough to reconstruct 360 scenes during inference time, taking 395 ms. Our method allows, for example, rendering third-person images and birds-eye views. Code, and more results are available at our project page.
Figure 1: Qualitative results of 6Img-to-3D. During inference time, 6Img-to-3D takes six surround vehicle RGB images with a small overlap as input and returns a parameterized triplane from which arbitrary novel views can be rendered.
@@ 1 Introduction
Inferring the appearance and geometry of large-scale outdoor scenes from a few camera inputs is a challenging, unsolved problem. The problem is characterized by the complexity inherent in outdoor scenes, namely a vast spatial extent, diverse object textures, and ambiguity of the 3D geometry due to occlusions. Robotics and autonomous driving systems both require methods that process vision-centric first-person views from a fixed set of cameras as input and derive suitable control commands from these [83]. The ability to instantaneously perform 2D-to-3D scene reconstructions would be useful within such methods and thus have broad applicability within robotics [2, 64, 92] and autonomous driving [16, 113, 9] domains. Quickly generating unobstructed bird’s-eye views of challenging parking scenarios or of large vehicles for teleoperation could significantly aid human drivers. For applying one-shot images to high-fidelity 3D techniques in vehicles or robots, additional factors, such as the inference speed, the quality of the inferred representation, the scalability of the approach, and the number of camera parameters required for the technique to work, are equally important. For those safety-critical domains, avoiding overly compressing the 3D scene or altering scene details by adding or omitting road obstacles is crucial.
**footnotetext: These authors contributed equally to this work.
Despite recent progress in 3D reconstruction from a few or single 2D images, many methods are either limited to single objects [29, 78, 51, 50, 99] or to generative indoor view synthesis [13]. Existing few-shot methods developed for single objects are often not easily transferable to large-scale or unbounded outdoor settings. This is because outdoor settings are more diverse due to their varying structural composition, occlusions, and vast extent. Another limitation is that many methods require a high degree of overlap between the input images, such as found in spherically arranged and inward-facing camera setups focused on a single object in the center [58, 105]. This is opposed to the outward-facing camera setup essential for autonomous driving where the camera overlap is minimal [6, 10, 76]. The camera setups are contrasted in Fig. 2. Traditional methods in autonomous driving have shown promising results with such camera setups but lack visual fidelity since they only predict semantic occupancy [32, 84, 90] and commonly rely on additional sensors such as LiDAR.
@@
Problem Statement. Our work addresses the challenge of designing, training, and evaluating a novel single-shot model for novel view synthesis able to render birds-eye and third-person perspectives of unbounded traffic scenes solely from six outward-facing ego vehicle images with minimal overlap. The relative camera poses are fixed and known. For training and evaluation, a synthetic dataset is used due to the absence of a suited real-world dataset.
Given the above limitations and the problem statement, we present 6Img-to-3D, a transformer-based decoder-renderer method for single-shot image-to-3D reconstruction. 6Img-to-3D is fully differentiable and end-to-end trained on a single GPU. It simultaneously takes six outward-facing images (6Img-) as inputs and generates a parameterized triplane (to-3D) from which novel scene views can be rendered. Unlike previous methods, we do not focus on single objects but instead on complex driving scenarios characterized by a high-depth complexity recorded only by sparse multi-view outward-facing RGB cameras. Furthermore, unlike other methods, we do not use a multi- or few-shot scheme where an initial reconstruction is tested and improved iteratively. Instead, we output the 3D scene in a single shot without iterative model refinement. To this end, we utilize pre-trained ResNet features, cross- and self-attention mechanisms for parameterizing the triplane, and image feature projection to condition the renderer. Our method is trained using roughly 11.4K input and 190K supervising images from 1900 scenes. We do not require depth or LiDAR information. Due to the absence of suitable datasets, including a large number of non-ego vehicle views needed during training, we have to rely on synthetic data from Carla [20].
@@
    Novel Architecture. We present 6Img-to-3D, a novel single-shot, few-image novel view synthesis pipeline for large-scale outdoor environments. Our method outputs a parameterized triplane of the 3D scene from which arbitrary viewpoints can be rendered. We combine triplane-based differentiable volume rendering, renderer conditioning on projected image features, self- and custom cross-attention mechanisms, and an LPIPS loss.
@@
    Visual Fidelity. We demonstrate that our method is, unlike other approaches, able to faithfully reconstruct unobserved scene parts. Our method outperforms existing few-image methods such as SplatterImage both in terms of depth and visual fidelity.
@@
    Hardware Efficiency. Unlike other large-scale 2D-to-3D methods, training 6Img-to-3D requires a single 42GB GPU and for inference only 10.5GB. From six-image-to-triplane our architecture only takes 395ms.
@@ 2 Related Work
3D Representations. Traditional methods for representing 3D scenes include voxels [71, 26, 93, 91], point clouds [21, 107, 102, 1, 108], signed distance field (SDF) [87] and polygon meshes [35, 86, 61, 97, 53]. Recently, numerous new implicit and explicit representations have emerged for learning 3D scenes [82, 95, 25, 38]. Implicit neural fields such as Neural Radiance Fields (NeRFs) model the surrounding appearance and geometry using a continuously queriable function updated through differentiable volumetric rendering [62, 58, 54, 79, 25], either utilize a single [17, 58, 111] or multiple [69, 67] multilayer perceptrons (MLPs). However, these models are slow to train, data-intensive, and expensive to render since querying the implicit neural field is computationally demanding. Explicit approaches such as Plenoctrees [104], Plenoxels [22], and NSVF [49] employ trilinear interpolation within a 3D grid to represent the scene. This significantly improves rendering and optimization time but comes at the expense of limited scalability due to the curse of dimensionality when increasing the resolution or scene size. 3D Gaussian Splats [38] as another explicit approach do not face those downsides since they do not require empty space to be explicitly modeled. Hybrid models [75, 60] adopt a voxel-like grid as an intermediary representation. Some models further refine this process by decomposing the voxel grid into three or more orthogonal planes called triplanes or K-Planes [15, 12, 5, 23]. This enables the computation of features for each point in space by aggregating the feature values of projected 3D points onto each plane, transitioning from an N-cube voxel grid to a 3 $\times$ N-squared plane representation.
@@
Few-Image to 3D Representation. Single-scene 3D reconstruction methods require many images to generalize towards novel views since the reconstruction problem would be underconstrained otherwise [58]. Hence, few-image to 3D representation methods incorporate additional global or local information to further regularize and constrain the problem. Global regularization methods rely on additional model priors, either via applying further model regularization [100, 63] or by utilizing pre-trained image models to provide an extra guidance signal to inform the optimization of the single-scene [19, 66, 50, 57, 72, 96, 99, 47, 65]. Local regularization methods train a cross-scene multi-perspective aggregator. Such models include PixelNeRF [105], IBRNet [88], MVSNeRF [14], VolRecon [70] and others [85, 52, 34, 45]. They retrieve image features via view projection and aggregate resulting features to obtain novel views. We also apply pixel feature conditioning, but unlike our method, many of the mentioned local regularization methods rely on a large overlap between input images ensured by using spherical inward-facing cameras. Another category of methods does not train the representation on a per-scene basis but instead trains a meta-network that outputs the discrete parameterized scene representation. Models such as [74, 24, 27, 3, 89] use diffusion to generate 3D representations of scenes. Many of those models use the triplane representation [74, 24, 27, 3, 8], whereas some directly generate in the image-space [89] or can be conditioned on input images [27, 89, 3]. The recent single-shot method SplatterImage [77] parameterizes several pixel-aligned Gaussians as scene representation. While being fast to render, this method struggles with occlusions and unobserved scene parts. The method that most closely resembles our work is LRM [29], which utilizes cross- and self-attention to parameterize a triplane using a few object images. However, LRM does not handle unbound scenes, is trained on 128 NVIDIA A100, and has seven times more parameters than our method.
@@
Large-Scale 3D Scene Representations. Extending NeRFs to unbounded scenes requires spatially contracting the scene, which is usually done proportionally to disparity as introduced in MipNeRF-360 [4] or using an inverted sphere parametrization as in NeRF++ [110]. Architectures modeling outdoor driving scenes are BlockNeRF [80], DisCoScene [98], NeuralField-LDM [39] and Neural ground planes [73]. Those are, however, not dealing with sparse input views. Neo360 [33] utilizes local image features to infer an image-conditional triplanar representation, whereby the model dissociates foreground from background scene parts. Unlike our method, Neo360 is trained on 8 A100 GPUs and focuses on inward-facing camera perspectives with large view overlap as input to the model. Large-scale scene generation methods such as Infinite Nature [48, 42] and others often produce additional images in an autoregressive fashion without the ability to input camera coordinates for novel view synthesis control [41, 94, 44, 18, 11]. In autonomous driving, established methods create a bird’s-eye-view (BEV) semantic representation of a scene using either Inverse Perspective Mapping (IPM) [68] or attention mechanisms [101, 43]. TPVFormer [31], a work that our architecture is partially based on, maps six outward-facing vehicle 2D images onto three orthogonal planes using self- and cross-deformable attention mechanisms to obtain semantic occupancy predictions. The method is, however, supervised with depth information from a LiDAR sensor and, unlike our method, only predicts semantic occupancy.
@@
Given six outward-facing images $\bm{I}_{ego}$, their camera extrinsic (relative pose) matrices $\bm{M}$ and associated camera intrinsic matrices $\bm{K}$ with appropriate dimensions, the goal is to reconstruct the surrounding 3D occupancy and appearance, that is the scene $\bm{S}$. Since the values of the scene $\bm{S}$ are not known, we use ${N}$ spherical multi-view images $\bm{I}_{sphere}$ with their associated extrinsic and extrinsic matrices during training as a supervision signal to estimate the scene $\bm{\hat{S}}$. The proposed architecture for lifting $\bm{I}_{ego}$ consists of an image-to-triplane encoder. After a forward pass through this architecture, the discrete triplane contains the approximation of the scene $\bm{\hat{S}}$ from which novel views $\bm{I}_{i}$ can be rendered. The following sections describe the different parts of the pipeline visualized in Fig. 3 and the training objectives.
@@ 3.3 Image-to-Triplane Encoder
$\begin{split}\textbf{CrossAttn}(\bm{T_{k}})=\text{DeformAttn}(\bm{T_{k}},\bm{p^{F}_{k}},\bm{F_{I}}),\\ k\in\{HW,HZ,WZ\}\ \end{split}$
@@
$\begin{split}\textbf{SelfAttn}(\bm{T_{k}})=\text{DeformAttn}(\bm{T_{k}},\bm{p^{T}_{k}},\mathbf{T}),\\ k\in\{HW,HZ,WZ\}\;.\end{split}$
@@ 3.5 Training Objectives
$\begin{split}\mathcal{L}=\mathcal{L}_{color}(\bm{C}(\bm{r}),\hat{\bm{C}}(\bm{r}))+\lambda_{t}\mathcal{L}_{TV}(\bm{T})\\ +\lambda_{d}\mathcal{L}_{dist}(\bm{s},\bm{w})+\lambda_{l}\mathcal{L}_{\text{LPIPS}}(\bm{x},\hat{\bm{x}})\;.\end{split}$ (3)
@@
$\begin{split}\mathcal{L}_{dist}(\bm{s},\bm{w})=\sum_{i,j}w_{i}w_{j}\bigg{|}\frac{s_{i}+s_{i+1}}{2}-\frac{s_{j}+s_{j+q}}{2}\bigg{|}\\ +\frac{1}{3}\sum_{i}w_{i}^{2}(s_{i+1}-s_{i})\;,\end{split}$ (6)
where $\bm{s}$ is a set of ray distances and $\bm{w}$ are the volume rendering weights parameterizing each ray, following [4].
@@ 4.1 Dataset
We use a synthetic dataset containing both ego and non-ego vehicle views few-view image novel view synthesis in an autonomous driving setting. Concerning novel view synthesis, other approaches also rely on synthetic data [33]. The dataset contains 2000 single-timestep complex outdoor driving scenes, each offering six outward-facing vehicle images and 100 spherical images for supervision. The dataset also consists of five multi-timestep scenes, each with 200 consecutive driving steps containing six outward-facing vehicle cameras, one bird’s-eye view, and one camera following the vehicle from behind. We only use the multi-timestep data for visualization. This results in 220K individual RGB images, each with their associated intrinsic camera matrix and pose. The dataset represents various driving scenes, vehicle types, pedestrians, and lighting conditions. To show the generalizability of our method, we use CARLA [20] Town 1, Town 3 to 7, and 10 for generating training data, resulting in 1900 training scenes, and left the 100 scenes from Town 2 for testing. For further details, we refer to the publication of the dataset [36].